# Physics of droplet regulation in biological cells

### v1

@@ 5.2.1  Stability of homogeneous states
where we used the exchange chemical potential ${\bar{\mu}=\nu\delta F/\delta\phi}$ corresponding from the free energy $F$ given in equation (2.22). The stationary state is unstable if perturbations grow, i.e., if $\omega(\boldsymbol{q})>0$. Without chemical reactions ($s^{\prime}(\phi_{0})=0$), this can only be the case if $f^{\prime\prime}(\phi_{0})<0$, which is the condition for the spinodal instability described in section 2.1.3. In this case, the fastest growing mode has a length scale $\ell_{\mathrm{max}}=2\pi/q_{\mathrm{max}}=\pi[-8\kappa/f^{\prime\prime}(\phi_{0})]^{1/2}$. As such, a homogeneous state would be expected to exhibit spinodal decomposition with characteristic length scale $\ell_{\mathrm{max}}$.
Figure 5.2: Stability of homogeneous states. Growth rate $\omega$ given by equation (5.7) as a function of wave number $|\boldsymbol{q}|$ of the perturbation for various reaction sensitivities $s^{\prime}(\phi_{0})$ with $\omega_{0}=\Lambda(\phi_{0})[f^{\prime\prime}(\phi_{0})]^{2}/4\kappa$ and $q_{\mathrm{max}}=(-f^{\prime\prime}(\phi_{0})/2\kappa)^{1/2}$.
Reactions affect the stability of the homogeneous state substantially; see figure 5.2. For $s^{\prime}(\phi_{0})>0$, the homogeneous state is always unstable because increasing $\phi$ further accelerates its production in an autocatalytic process. In contrast, the homogeneous state is completely stabilised for $s^{\prime}(\phi_{0})<k_{\mathrm{stab}}$, where $k_{\mathrm{stab}}=-\Lambda[f^{\prime\prime}(\phi_{0})]^{2}/4\kappa$, i.e., reactions that are faster than this rate completely suppress spinodal decomposition. Most interestingly, there also exists an intermediate regime, $k_{\mathrm{stab}}<s^{\prime}(\phi_{0})<0$, where only a compact band of wavelengths is unstable. In all cases, the most unstable mode has wavelength $l_{\mathrm{max}}$, and so expect initial spinodal decomposition to proceed as expected in the intermediate regime. In particular, droplets will then form and Ostwald ripening (described in section 2.2.4) may kick in. However, that fact that $\omega$ is negative for long wavelengths (small $\boldsymbol{q}$) in this intermediate regime suggests that the late stage dynamics might be different. To investigate this, we next analyse the dynamics of multiple droplets.
@@ 5.2.2  Effective droplet model
To understand the behaviour of chemically active droplets in more detail, we next discuss the effective droplet model introduced in section 2.2. The equilibration of the interfacial region is not affected significantly by the reactions since the interfacial volume is small. Instead, reactions affect the dynamical equation in the bulk phases, which after linearisation led to the diffusion equation (2.37). Repeating this linearization for equation (5.6), we find
@@
inside droplets ($n=\mathrm{in}$) and in the surrounding dilute phase ($n=\mathrm{out}$) with the familiar diffusivities ${D_{n}\approx\nu\Lambda(\phi^{(0)}_{n})f^{\prime\prime}(\phi^{(0)}_{n})}$. Here, we have defined the basal reaction flux $\Gamma_{n}=s(\phi^{(0)}_{n})$, which quantifies the production of droplet material at the equilibrium fraction $\phi^{(0)}_{n}$. We also introduce reaction rates ${k_{n}=-s^{\prime}(\phi^{(0)}_{n})}$, which specify how the production of droplet material changes when concentrations deviate from equilibrium.
The reaction-diffusion equation (5.8) is linear in $\phi$ and can be solved analytically for spherical droplets of radius $R$ [140]. The resulting concentration profiles exhibit characteristic length scales given by the reaction-diffusion lengths $\ell_{n}=\sqrt{D_{n}/k_{n}}$, which are defined for $k_{n}>0$, i.e., when reactions are stabilising. In particular, the concentration profile is generally heterogeneous inside and outside the droplet, implying diffusive fluxes $\boldsymbol{j}_{n}$ on both sides of the interface, which affect the droplet growth speed given by equation (2.36). In the limit of small droplets, $R\ll\ell_{\mathrm{in}}$, the diffusive flux inside the interface is balanced by the reaction in the entire droplet volume, resulting in the approximate growth rate [397]
@@
where $\phi_{\infty}=\phi^{(0)}_{\mathrm{out}}+\Gamma_{\mathrm{out}}/k_{\mathrm{out}}$ is the volume fraction far away from the droplet and the equilibrium fraction $\phi^{\mathrm{eq}}_{\mathrm{out}}$ is given by equation (2.32). Equation (5.9) differs in two crucial aspects from equation (2.39) describing passive droplets. First, the fraction $\phi_{\infty}$ is now a constant value controlled by the reactions in the dilute phase, in contrast to the passive case where it was determined by material conservation and depended on the size of all droplets. Indeed the dynamics of multiple droplets are independently described by equation (5.9), as long as the droplet separation is large enough compared to $\ell_{\mathrm{out}}$ such that they do not disturb their corresponding concentration profiles. Second, equation (5.9) contains a term related to the reactions inside the droplet, which could dominate growth over the surface tension effects captured by the other terms in the bracket.
@@ 5.2.3  Two classes of active droplets
Equation (5.9) suggests that the basal fluxes $\Gamma_{n}$ can affect the dynamics of chemically active droplets significantly. This is not surprising since they generally determine whether droplet material is produced ($\Gamma_{n}>0$) or degraded ($\Gamma_{n}<0$) inside the droplet ($n=\mathrm{in}$) and in its surrounding ($n=\mathrm{out}$). If droplet material is degraded (produced) everywhere, the system will end up in a dilute (dense) homogeneous state; see grey lines in figure 5.1. To have stable droplets, material thus needs to be produced in one phase while it is degraded in the other, leaving us with two distinct classes of active droplets: externally maintained droplet and internally maintained droplets. Externally maintained droplets degrade droplet material inside the droplet ($\Gamma_{\mathrm{in}}<0$), while it is produced outside the droplet ($\Gamma_{\mathrm{out}}>0$). A simple example of a possible reaction rate for this case is shown in figure 5.1a. On the other hand, internally maintained droplets produce droplet material inside the droplet ($\Gamma_{\mathrm{in}}>0$), while it is degraded outside the droplet ($\Gamma_{\mathrm{out}}<0$). The green line in figure 5.1b shows that in this case there are stable states outside the blue binodal region, implying that a dilute and a dense homogeneous state can be stable. If the two stable points are instead inside the binodal region, we obtain a more complex version of externally maintained droplets (orange line), whereas just one homogeneous state survives if exactly one stable point is inside the binodal (grey lines). Taken together, this qualitative analysis suggests that there are two classes of chemically active droplets, which have very different behaviours as summarised in Table 5.1, and which we discuss separately in the following sections.
@@
Additional features Droplet division [139] Active shells [401] Self-organized drift [402] Stabilization by active cores [383] Centering of active cores [403]
@@ 5.3 Externally maintained droplets
Externally maintained droplets arise when the volume faction of the dilute phase away from the droplets is maintained at a volume fraction $\phi_{\infty}>\phi^{(0)}_{\mathrm{out}}$, where we recall that $\phi^{(0)}_{\mathrm{out}}$ is the volume fraction imposed at the interface by phase equilibrium. The associated supersaturation $\mathcal{S}=\Gamma_{\mathrm{out}}/(k_{\mathrm{out}}\phi^{(0)}_{\mathrm{out}})$ results from a net production of droplet material in the dilute phase ($\Gamma_{\mathrm{out}}>0$), which is balanced by degradation of droplet material within the dense phase (${\Gamma_{\mathrm{in}}<0}$). Figure 5.3a demonstrates that these opposing reactions cause diffusive fluxes across the interface, transporting droplet material into the droplet in exchange for precursor material. This transfer of material governs the dynamic behaviour described by equation (5.9). For externally maintained droplets, this equation has two stationary solutions given by,
@@
Linear stability analysis of equation (5.9) reveals that the first of these solutions (equation (5.10a)) is unstable and thus corresponds to the critical radius, which must be exceeded for droplet nucleation to occur. The second term in equation (5.10a) indicates that the critical radius increases for stronger reactions (more negative $\Gamma_{\mathrm{in}}$) even when the supersaturation $\mathcal{S}$ is kept fixed. We analyse in detail the suppression of droplet nucleation by reactions in section 5.3.3 below. The second stationary solution, (equation (5.10b)), is stable, implying that externally maintained droplets prefer to attain a fixed finite size.
@@ 5.3.1  Droplet size control
The expression for the stable stationary radius $R_{*}$ derived above relies on the assumption that a single active droplet is surrounded by a large dilute phase at an average fraction $\phi_{0}$. In this case, droplets that are larger than $R_{*}$ will shrink towards this size, since they will lose more droplet material via internal degradation reactions than they will gain via the diffusive influx from the dilute phase. Equation (5.10b) shows that this size control ceases in passive systems since $R_{*}\rightarrow\infty$ for $\Gamma_{\mathrm{in}}\rightarrow 0$. The fact that isolated droplets attain a fixed size suggests that the surface tension driven coarsening process of Ostwald ripening (described in section 2.2.4) is suppressed. A more detailed calculation shows that multiple externally maintained droplets ($\Gamma_{\mathrm{in}}<0$) can stably coexist if they are larger than the threshold radius [399]
@@
Whether multiple externally maintained droplets coexist or a single droplet dominates the system not only depends on the physical parameters, such as reaction rates, but also on initial conditions [399, 406]. Because a range of droplet sizes can be stable, the precise initialisation protocol determines the final state. In particular, starting with a single small droplet, there exists two qualitatively different regimes [406]. Either, the droplet grows until it reaches a fixed radius comparable to $R_{*}$, or the droplet grows extensively, occupying a fixed fraction of the system. In the latter case, the droplet radius eventually exceeds the reaction-diffusion length $\ell_{\mathrm{in}}$, so that the approximation leading to equation (5.9) breaks down and the solution given in equation (5.10b) no longer applies. Essentially, such large active droplets have chemical fluxes restricted to a region of size $\ell_{n}$ around their interfaces, permitting extensive growth overall. Taken together, this implies that externally maintained droplets can attain multiple stable states, but the details are currently not well understood.
@@ 5.3.2  Electrostatic analogy
Some behaviours of externally maintained droplets can be discussed intuitively in the special case of linearised reactions, ${s(\phi)=k(\phi_{0}-\phi)}$, where $k$ determines the reaction rate, and $\phi_{0}$ denotes the fraction in the dilute phase. In this case, the dynamics of the entire system obey the Cahn-Hilliard-Onoo equation [407]
@@
where $\Lambda$ is a constant mobility. Phase-separated structures emerging from these dynamics show size selection [408, 409, 410, 411].
Although equation (5.12) describes an active system, it can be written in the form of gradient dynamics, $\partial_{t}\phi=\Lambda\nabla^{2}\delta\tilde{F}/\delta\phi$, with a modified free energy
@@
with appropriate boundary conditions [412]. In analogy to the electrostatic interactions discussed in section 3.1, $\varphi$ can thus be interpreted as the electrostatic potential, $\phi(\boldsymbol{r})-\phi_{0}$ plays the role of the charge density $\varrho(\boldsymbol{r})$, and $\Lambda/k$ is equivalent to the electric permittivity $\varepsilon$. Figure 5.4 shows how the effect of the chemical reactions can thus be interpreted as a charge cloud surrounding oppositely charged droplets. In general, large droplets will accrue a lot of charge, increasing the energy corresponding to the equivalent electrostatic interaction, such that further growth is halted.
The major advantage of this electrostatic analogy is that states that minimise $\tilde{F}$ are necessarily stationary states of equation (5.12), and we can use equilibrium thermodynamics to analyse their behaviour. This method has been successful in analysing length scale selection [409, 410, 411, 412, 413], in analogy with the seminal Ohta-Kawasaki model [414, 415] and other pattern-forming models based on long-range repulsion [416]. Consequently, patterns generally emerge in systems that combine short-range attraction by phase separation with long-range repulsion due to chemical reactions [417, 418, 419, 420, 421, 422, 423, 424, 425, 426].
@@ 5.3.3  Suppressed nucleation
The electrostatic analogy introduced above also explains why droplet nucleation is suppressed by reactions [398]. The nucleation kinetics in this system are well-described by the classical nucleation theory described in section 2.2.2 [427]. Following the electrostatic analogy, the energy penalty $\Delta\!F$ associated with forming a small nucleus of radius $R$ approximately reads [413]
@@
where the first two terms are equivalent to the passive case given by equation (2.33) and the last term captures the influence of reactions in the limit $\ell\gg R$. If droplet formation is favourable ($\Delta\!f>0$) and reactions are weak, $\Delta\!F$ exhibits a maximum at $R\approx R_{\mathrm{crit}}$ given by equation (5.10a), which reads
@@
to linear order in $k$. The pre-factor here is identical to the passive energy barrier given by equation (2.34). Reactions thus increase $\Delta\!F_{\mathrm{nucl}}$, implying a smaller nucleation rate $J\propto\exp(-\Delta\!F_{\mathrm{nucl}})$; see equation (2.35). Since chemical reactions can be interpreted as a charge separation, this suppression of nucleation follows from the fact that forming a charged nucleus is energetically unfavourable compared to forming a neutral nucleus. Another interpretation is the fact that the chemical reactions want to maintain a fixed concentration locally and thus dampen fluctuations originating from diffusive fluxes, making the formation of a critical nucleus less likely.
Chemical reactions also affect heterogeneous droplet nucleation at surfaces [413]. In section 4.1.4, we discuss how surfaces promote nucleation, and this behaviour is retained when reactions describing externally maintained droplets are added [413]. However, sufficiently large reaction rates will suppress nucleation, as for the homogeneous case. Moreover, reactions may cause deformation of sessile droplets at surfaces (see figure 5.4b), as discussed in more detail in section 5.5.1 below. The intertwined interactions of surfaces and reactions thus provide flexible control over droplet nucleation.

### v2

@@ 5.2.1  Stability of homogeneous states
Figure 5.2: Stability of homogeneous states. Growth rate $\omega$ given by equation (5.8) as a function of wave number $|\boldsymbol{q}|$ for various reaction sensitivities $s^{\prime}(\phi_{0})$ with ${\omega_{0}=\kappa^{-1}\Lambda(\phi_{0})[f^{\prime\prime}(\phi_{0})]^{2}}$ and $q_{\mathrm{max}}=(-f^{\prime\prime}(\phi_{0})/2\kappa)^{1/2}$.
@@
where we have used the exchange chemical potential ${\bar{\mu}=\nu\delta F/\delta\phi}$ corresponding to the free energy $F$ given in equation (2.22). The stationary state is unstable if perturbations grow, i.e., if $\omega(\boldsymbol{q})>0$ for any value of  $\boldsymbol{q}$. Without chemical reactions, $s^{\prime}(\phi_{0})=0$, this is the case if $f^{\prime\prime}(\phi_{0})<0$, corresponding to the spinodal instability described in section 2.1.3.
The influence of chemical reactions on the stability of the homogeneous system is more subtle (Figure 5.2). For autocatalytic reactions where $s^{\prime}(\phi_{0})>0$, the homogeneous state is always unstable since increasing $\phi$ further accelerates its production, destabilizing the overall composition, $\omega(\boldsymbol{q}=0)>0$. In contrast, the overall composition is stable $\omega(\boldsymbol{q}=0)<0$ for auto-inhibitory reactions where $s^{\prime}(\phi_{0})<0$. In this case, the fastest growing mode corresponds to a length scale $\ell_{\mathrm{max}}=2\pi/q_{\mathrm{max}}=\pi[-8\kappa/f^{\prime\prime}(\phi_{0})]^{1/2}$ and grows at rate $\omega_{\mathrm{max}}=s^{\prime}(\phi_{0})+\Lambda[f^{\prime\prime}(\phi_{0})]^{2}/4\kappa$. Consequently, auto-inhibitory reactions only exhibit instabilities in the spinodal region where $f^{\prime\prime}(\phi_{0})<0$. However, even in the spinodal region, the homogeneous state can be completely stabilised if the reactions are sufficiently strong, $s^{\prime}(\phi_{0})<k_{\mathrm{stab}}$, where $k_{\mathrm{stab}}=-\Lambda[f^{\prime\prime}(\phi_{0})]^{2}/4\kappa$. For weaker reactions, $k_{\mathrm{stab}}<s^{\prime}(\phi_{0})<0$, only a compact band of wavelengths is unstable (Figure 5.2). If $f^{\prime\prime}(\phi_{0})<0$, the length scale $\ell_{\mathrm{max}}$ of the most unstable mode is independent of $s^{\prime}(\phi_{0})$, so we expect standard spinodal decomposition to initially develop in the intermediate regime. In particular, droplets will form and Ostwald ripening (see section 2.2.4) may drive coarsening. However, the fact that $\omega$ is negative for long wavelengths (small $\boldsymbol{q}$) in this intermediate regime suggests that late stage dynamics may deviate from the classical ripening behaviour. To investigate this possibility, we next analyse the dynamics of multiple droplets.
@@ 5.2.2  Effective droplet model
To understand the dynamics of chemically active droplets in more detail, we next discuss the effective droplet model introduced in section 2.2. The equilibration of the interfacial region is not affected significantly by reactions since the interfacial volume is typically small. Instead, reactions affect the dynamical equation in the bulk phases, which after linearisation led to the diffusion equation (2.37). Repeating this linearisation for equation (5.7), we find
@@
both inside droplets ($n=\mathrm{in}$) and in the surrounding dilute phase ($n=\mathrm{out}$), with the familiar diffusivities ${D_{n}\approx\nu\Lambda(\phi^{(0)}_{n})f^{\prime\prime}(\phi^{(0)}_{n})}$. Here, we also define the basal reaction flux $\Gamma_{n}=s(\phi^{(0)}_{n})$, which quantifies the production of droplet material at the equilibrium fraction $\phi^{(0)}_{n}$. We also introduce reaction rates ${k_{n}=-s^{\prime}(\phi^{(0)}_{n})}$, which specify how the production of droplet material changes when concentrations deviate from equilibrium.
The reaction–diffusion equation (5.9) is linear in $\phi$ and can be solved analytically for spherical droplets of radius $R$ [140]. The resulting concentration profiles display spatial variations over characteristic length scales given by the reaction–diffusion lengths $\ell_{n}=\sqrt{D_{n}/k_{n}}$, which are defined for $k_{n}>0$ (stabilising reactions). In general, these variations drive diffusive fluxes $\boldsymbol{j}_{n}$ on both sides of the interface, which affect the droplet growth speed given by equation (2.36). In the limit of small droplets, $R\ll\ell_{\mathrm{in}}$, the diffusive flux across the interface is balanced by the reaction in the entire droplet volume, resulting in the approximate growth rate [406]
@@
where $\phi_{\infty}=\phi^{(0)}_{\mathrm{out}}+\Gamma_{\mathrm{out}}/k_{\mathrm{out}}$ is the volume fraction far away from the droplet and the equilibrium fraction $\phi^{\mathrm{eq}}_{\mathrm{out}}$ is given by equation (2.32). Equation (5.10) differs in two crucial aspects from equation (2.39) describing passive droplets. First, the fraction $\phi_{\infty}$ is now a constant value controlled by the reactions in the dilute phase, in contrast to the passive case where it was determined by material conservation and depended on the size of all droplets. Indeed the dynamics of multiple droplets are independently described by equation (5.10), as long as the droplet separation is large enough compared to $\ell_{\mathrm{out}}$ such that they do not disturb their corresponding concentration profiles. Second, equation (5.10) contains a term related to the reactions inside the droplet, which can dominate growth over the surface tension effects captured by the other terms in this equation.
@@ 5.2.3  Two classes of active droplets
Equation (5.10) suggests that the basal fluxes $\Gamma_{n}$ can affect the dynamics of chemically active droplets significantly. This is not surprising since these fluxes determine whether droplet material is produced ($\Gamma_{n}>0$) or degraded ($\Gamma_{n}<0$) inside the droplet ($n=\mathrm{in}$) and in its surrounding ($n=\mathrm{out}$). If droplet material is degraded (produced) everywhere, the system will end up in a dilute (dense) homogeneous state (see grey lines in figure 5.1). To have stable droplets, material thus needs to be produced in one phase while it is degraded in the other, leaving two distinct classes of active droplets: externally maintained droplets and internally maintained droplets. Externally maintained droplets degrade droplet material inside the droplet ($\Gamma_{\mathrm{in}}<0$), while it is produced outside the droplet ($\Gamma_{\mathrm{out}}>0$). A simple example of a possible reaction rate for this case is shown in figure 5.1a. On the other hand, internally maintained droplets produce droplet material inside the droplet ($\Gamma_{\mathrm{in}}>0$), while it is degraded outside the droplet ($\Gamma_{\mathrm{out}}<0$). The green line in figure 5.1b shows that in this case there are stable states outside the blue binodal region, implying that a dilute and a dense homogeneous state can be stable. If the two stable points are instead inside the binodal region, we obtain a more complex version of externally maintained droplets (orange line), whereas just one homogeneous state survives if exactly one stable point is inside the binodal region (grey lines). Taken together, this qualitative analysis suggests that there are two classes of chemically active droplets, which have very different behaviours as summarised in Table 5.1, and which we discuss separately in the following sections.
@@
Additional features
Droplet division [139]
Active shells [410]
Self-organised drift [411]
Stabilisation by active cores [392]
Centring of active cores [412]
@@ 5.3 Externally maintained droplets
Externally maintained droplets form when the volume fraction of the dilute phase away from droplets is maintained at a volume fraction $\phi_{\infty}>\phi^{(0)}_{\mathrm{out}}$, where we recall that $\phi^{(0)}_{\mathrm{out}}$ is the volume fraction imposed at the interface by phase equilibrium. The associated supersaturation $\mathcal{S}=\Gamma_{\mathrm{out}}/(k_{\mathrm{out}}\phi^{(0)}_{\mathrm{out}})$ results from a net production of droplet material in the dilute phase ($\Gamma_{\mathrm{out}}>0$), which is balanced by degradation of droplet material within the dense phase (${\Gamma_{\mathrm{in}}<0}$). Figure 5.3a demonstrates that these opposing reactions cause diffusive fluxes across the interface, transporting droplet material into the droplet in exchange for precursor material. This transfer of material governs the dynamic behaviour described by equation (5.10). For externally maintained droplets, this equation has two stationary solutions given by,
@@
Linear stability analysis of equation (5.10) reveals that the first of these solutions (equation 5.11a) is unstable and thus corresponds to the critical radius that must be exceeded for droplet nucleation to occur. The second term in equation (5.11a) indicates that the critical radius increases for stronger reactions (more negative $\Gamma_{\mathrm{in}}$) even when the supersaturation $\mathcal{S}$ is kept fixed. We analyse in detail the suppression of droplet nucleation by reactions in section 5.3.3 below. The second stationary solution (equation 5.11b) is stable, implying that externally maintained droplets prefer to attain a fixed finite size.
@@ 5.3.1  Droplet size control
The expression for the stable stationary radius $R_{*}$ derived above relies on the assumption that a single active droplet is surrounded by a large dilute phase at an average fraction $\phi_{0}$. In this case, droplets that are larger than $R_{*}$ will shrink, since they lose more droplet material via internal degradation reactions than they gain via the diffusive influx from the dilute phase. Equation (5.11b) shows that this size control ceases in passive systems since $R_{*}\rightarrow\infty$ for $\Gamma_{\mathrm{in}}\rightarrow 0$. The fact that isolated droplets attain a fixed size suggests that droplet coarsening via Ostwald ripening is suppressed. A more detailed calculation shows that multiple externally maintained droplets ($\Gamma_{\mathrm{in}}<0$) can stably coexist if they are larger than the threshold radius [408]
@@
Whether multiple externally maintained droplets coexist or a single droplet dominates the system not only depends on physical parameters such as reaction rates, but also on initial conditions [408, 415]. Since a range of droplet sizes can be stable, the precise initialisation protocol determines the final state. In particular, starting with a single small droplet, there exists two qualitatively different regimes [415]: either the droplet grows until it reaches a fixed radius comparable to $R_{*}$, or it grows extensively, occupying a fixed fraction of the system. In the latter case, the droplet radius eventually exceeds the reaction-diffusion length $\ell_{\mathrm{in}}$, so that the approximation leading to equation (5.10) breaks down and the solution given in equation (5.11b) no longer applies. Essentially, these large active droplets have chemical fluxes restricted to a region of size $\ell_{n}$ around their interfaces, permitting extensive growth overall. Taken together, these arguments imply that externally maintained droplets can attain multiple stable states, although the details of this situation are currently not well understood.
@@ 5.3.2  Electrostatic analogy
Some behaviours of externally maintained droplets can be discussed intuitively in the special case of linearised reactions, ${s(\phi)=k(\phi_{0}-\phi)}$, where $k$ determines the reaction rate, and $\phi_{0}$ denotes the fraction in the dilute phase. In this case, the dynamics of the entire system obey the Cahn–Hilliard–Onoo equation [416]
@@
where $\Lambda$ is a constant mobility. Phase-separated structures emerging from these dynamics generically show size selection [417, 418, 419, 420].
Although equation (5.13) describes an active system, it can also be written in the form of gradient dynamics, $\partial_{t}\phi=\Lambda\nabla^{2}\delta\tilde{F}/\delta\phi$, with a modified free energy
@@
with appropriate boundary conditions [421]. In analogy to the electrostatic interactions discussed in section 3.1, $\varphi$ can thus be interpreted as the electrostatic potential, $\phi(\boldsymbol{r})-\phi_{0}$ plays the role of the charge density $\varrho(\boldsymbol{r})$, and $\Lambda/k$ is equivalent to the electric permittivity $\varepsilon$. Figure 5.4 shows how the effect of the chemical reactions can thus be interpreted as a charge cloud surrounding oppositely charged droplets. In general, large droplets will accrue a large amount of charge, increasing the energy corresponding to the equivalent electrostatic interaction, such that further growth is halted.
A major strength of this electrostatic analogy is that states that minimise $\tilde{F}$ are necessarily stationary states of equation (5.13), and so we can use equilibrium thermodynamics to analyse their behaviour. This method has previously been successful in analysing length scale selection [418, 419, 420, 421, 422], in analogy with the seminal Ohta–Kawasaki model [423, 424] and other pattern-forming models based on long-range repulsion [425]. In general, patterns tend to emerge in systems that combine short-range attraction by phase separation with long-range repulsion due to chemical reactions [426, 427, 428, 429, 430, 431, 432, 433, 434, 435].
@@ 5.3.3  Suppressed nucleation
The electrostatic analogy introduced above also explains why droplet nucleation is suppressed by reactions [407]. Nucleation kinetics in this system are well described by the classical nucleation theory described in section 2.2.2 [436]. Following the electrostatic analogy, the energy penalty $\Delta\!F$ associated with forming a small nucleus of radius $R$ approximately reads [422]
@@
where the first two terms are equivalent to the passive case given by equation (2.33) and the last term captures the influence of reactions in the limit $\ell_{\mathrm{in}},\ell_{\mathrm{out}}\gg R$. If droplet formation is favourable ($\Delta\!f>0$) and reactions are weak, $\Delta\!F$ exhibits a maximum at $R\approx R_{\mathrm{crit}}$ given by equation (5.11a), which reads
@@
to linear order in $k$. The pre-factor here is identical to the passive energy barrier given by equation (2.34). Reactions thus increase $\Delta\!F_{\mathrm{nucl}}$, implying a smaller nucleation rate $J\propto\exp(-\Delta\!F_{\mathrm{nucl}})$; see equation (2.35). Since chemical reactions can be interpreted as a separation of charge, this suppression of nucleation follows from the fact that forming a charged nucleus is energetically unfavourable compared to forming a neutral nucleus. This process can also be understood from the fact that chemical reactions want to maintain a fixed concentration locally and thus dampen fluctuations originating from diffusive fluxes, making the formation of a critical nucleus less likely.
Chemical reactions also affect heterogeneous droplet nucleation at surfaces [422]. In section 4.1.4, we discuss how surfaces promote nucleation, and this behaviour is retained when reactions describing externally maintained droplets are added [422]. However, sufficiently large reaction rates will still suppress nucleation, as for the homogeneous case. Further, reactions may cause deformation of sessile droplets at surfaces (see figure 5.4b), as discussed in more detail in section 5.5.1 below. The interactions of droplets with both surfaces and active reactions thus provide flexible control over droplet nucleation.